# Video Diffusion Transformers Quantization

### v2

@@ 4.1 Setup
Video Generation Evaluation Settings: We apply -Quant to HunyuanVideo (Kong et al., 2024) and generate videos using a 50-step flow matching scheduler (Lipman et al., 2023) with embedded CFG (Ho & Salimans, 2022) scale 6.0 and flow shift factor 7.0. We also apply -Quant to Wan2.1 (Wan et al., 2025), detailed configurations and results are shown in the supplementary file. For evaluation, we adpot the benchmark suite provided by VBench (Huang et al., 2024) to comprehensively assess the quality of the generated videos. Specifically, we evaluate the model across eight major dimensions (Ren et al., 2024) that reflect key aspects of video generation. These metrics are designed to align closely with human perception, ensuring a reliable and standardized assessment of model performance. We compare -Quant with MinMax (Jacob et al., 2018b), SmoothQuant (Xiao et al., 2023), Quarot (Ashkboos et al., 2024), and the current SOTA video PTQ method ViDiT-Q (Zhao et al., 2025). All experiments are conducted on a single RTX4090 GPU.
@@ 4.2 Main Results
VBench Quantitative Comparison. As shown in Tab. 1, -Quant achieves significant improvements over several state-of-the-art quantization methods under two configurations. (1) High Precision Regime: When weights are set to 4 bits, existing methods require 8-bit activations to maintain basic functionality, while our configuration demonstrates superior performance with lower activation precision. Notably, our W4A6 mixed-precision configuration nearly matches the BF16 HunyuanVideo model while significantly outperforming all W4A8 baselines across most metrics. (2) Low Precision Challenge: In the extremely challenging W4A4 setting where all baseline methods either fail completely or suffer severe degradation, -Quant maintains remarkable stability. For example, we achieve 58.94 Aesthetic Quality and 60.38 Imaging Quality, outperforming the best W4A4 baseline by +10.53 in Aesthetic Quality. Besides, our method preserves temporal dynamics while maintaining high motion smoothness, whereas prior works either degrade visually or fail to generate coherent motion.
Qualitative Comparison. In W4A8 settings, while baseline methods preserve only coarse outlines, they still suffer from significant loss of fine details. As illustrated in Fig. 5, ViDiT-Q’s W4A8 outputs show washed-out textures in elements like launch towers. In contrast, our method consistently recovers these subtle features, demonstrating robust performance. The advantage becomes more evident in more challenging W4A4 settings, where conventional methods fail completely, generating either incoherent noise patterns or severely distorted outputs. Our method maintains remarkable visual coherence with marginal quality degradation compared to the BF16 baseline.
@@ 4.3 Ablation Study
BGR and ARQ. We validate the effectiveness of BGR and ARQ in Tab. 3. The full model with both BGR and ARQ achieves optimal performance across all metrics, demonstrating their synergistic effects. Removing only BGR leads to significant drops in VBench scores, indicating the critical role of progressively narrowing the weight quantization bound to preserve perceptual fidelity. Conversely, disabling ARQ components results in comparable declines, highlighting the importance of dynamically adjusting quantization scales during inference. Both BGR and ARQ maintain their effectiveness irrespective of -GBS.
| Method |
@@
| Methods |
| BGR | ARQ | ||||||||||||||||||
| ✓ | W4A6 | 58.15 | 58.68 | 98.04 | 98.49 | 98.21 | |||||||||||||
| ✓ | W4A6 | 57.85 | 57.72 | 98.23 | 97.86 | 98.10 | |||||||||||||
| ✓ | ✓ | W4A6 | 60.46 | 61.93 | 98.91 | 98.95 | 98.40 | ||||||||||||
| ✓ | W4A4 | 53.95 | 52.67 | 97.92 | 98.71 | 97.89 | |||||||||||||
| ✓ | W4A4 | 43.26 | 58.31 | 95.36 | 96.08 | 97.35 | |||||||||||||
| ✓ | ✓ | W4A4 | 59.57 | 58.93 | 98.67 | 99.00 | 98.47 | ||||||||||||
Comparison with other mixed-precision strategies. In the experiment, we set bit and bit in Eq. 11. Our method achieves an average bit width of 6 across 50 timesteps. We compare the proposed -Guided Bit Switching with other fixed-pattern mixed-precision strategies. Specifically, we compare with four representative approaches: (1) Static Temporal Partitioning (STP): allocate 4-bit precision for the first 25 timesteps and 8-bit precision for the subsequent 25 timesteps; (2) Inverse Temporal Partitioning (ITP): reverse the bitwidth allocation order of STP with 8-bit initially and 4-bit later; (3) Alternating Bitwidth Switching (ABS): dynamically alternate between 4-bit and 8-bit precision at each timestep; (4) Stochastic Bitwidth Allocation (SBA): randomly assign 4-bit or 8-bit precision to each timestep while maintaing the average bit number 25. As shown in Tab. 2, -Guided Bit Switching surpasses other static mixed-precision strategies in imaging quality. This error-driven dynamic decision mechanism effectively resolves the challenge of bit allocation in conventional mixed-precision strategies.
Threshold of -Guided Bit Switching. In our mixed-precision strategy, the threshold determines when to switch bit-width during inference. Empirically, a lower leads to more time steps using high-bit quantization, improving visual quality at the cost of increased bit-width. This presents a trade-off between performance and resource usage. Thus, we conduct a series of experiments with different thresholds. As shown in Fig. 6, our experiments demonstrate that as increases, the average bit-width decreases, leading to a gradual reduction in imaging quality. Notably, a key advantage of -GBS over existing methods is its smooth and continuous bit-width adaptation: when , the model consistently uses W4A8; when , it reduces to W4A4. Critically, for intermediate values, -GBS dynamically interpolates between 4-bit and 8-bit at the activation level, ensuring a continuous precision transition based on input characteristics. This smooth adaptation is in stark contrast to conventional approaches, which typically switch abruptly between discrete bit-widths, causing instability in output quality.
| Method |
| -Quant+TeaCache | W4A8 | 61.39 | 62.96 | 98.24 | 98.73 | 98.75 | 4.01 | |||||||||||||
@@ 4.4 Memory and Latency
-Quant achieves significant improvements in both memory efficiency and inference speed. As shown in Tab. LABEL:tab:storage, compared to the BF16 baseline (16/16 bit-width), our W4A8 configuration reduces memory usage by 3.68 while accelerating latency by 1.75. More aggressive quantization (W4A4) maintains the same memory savings but further boosts latency speedup to 2.12. Notably, W4A6 means roughly half the denoising steps use W4A8 and the other half use W4A4. This allows direct use of the widely adopted W4A4 and W4A8 GEMM kernels, eliminating the need to design dedicated mixed-precision kernels.
@@ 4.5 Integrate with Cache Mechanism
To demonstrate the orthogonal compatibility of -Quant with other compression paradigms, we integrate it with one of the SOTA cache methods TeaCache (Liu et al., 2024a). As illustrated in Tab. LABEL:tab:teacache, -Quant maintains its compression efficiency alongside TeaCache (Liu et al., 2024a), achieving additive speedup (up to 4.85) without degrading key video generation metrics. This validates the practicality of -Quant in resource constrained scenarios where multiple optimization dimensions could be jointly addressed.

### v3

@@ 4.1 Setup
Video Generation Evaluation Settings: We apply -Quant to HunyuanVideoΒ (Kong et al., 2024) and generate videos using a 50-step flow matching schedulerΒ (Lipman et al., 2023) with embedded CFGΒ (Ho and Salimans, 2022) scale 6.0 and flow shift factor 7.0. We also apply -Quant to Wan2.1Β (Wan et al., 2025), detailed configurations and results are shown in the supplementary file. For evaluation, we adpot the benchmark suite provided by VBenchΒ (Huang et al., 2024) to comprehensively assess the quality of the generated videos. Specifically, we evaluate the model across eight major dimensionsΒ (Ren et al., 2024) that reflect key aspects of video generation. These metrics are designed to align closely with human perception, ensuring a reliable and standardized assessment of model performance. We compare -Quant with MinMaxΒ (Jacob et al., 2018b), SmoothQuantΒ (Xiao et al., 2023), QuarotΒ (Ashkboos et al., 2024), and the current SOTA video PTQ method ViDiT-QΒ (Zhao et al., 2025). All experiments are conducted on a single RTX4090 GPU.
@@ 4.2 Main Results
VBench Quantitative Comparison.β As shown in Tab.Β 1, -Quant achieves significant improvements over several state-of-the-art quantization methods under two configurations. (1) High Precision Regime: When weights are set to 4 bits, existing methods require 8-bit activations to maintain basic functionality, while our configuration demonstrates superior performance with lower activation precision. Notably, our W4A6 mixed-precision configuration nearly matches the BF16 HunyuanVideo model while significantly outperforming all W4A8 baselines across most metrics. (2) Low Precision Challenge: In the extremely challenging W4A4 setting where all baseline methods either fail completely or suffer severe degradation, -Quant maintains remarkable stability. For example, we achieve 58.94 Aesthetic Quality and 60.38 Imaging Quality, outperforming the best W4A4 baseline by +10.53 in Aesthetic Quality. Besides, our method preserves temporal dynamics while maintaining high motion smoothness, whereas prior works either degrade visually or fail to generate coherent motion.
Qualitative Comparison.β In W4A8 settings, while baseline methods preserve only coarse outlines, they still suffer from significant loss of fine details. As illustrated in Fig.Β 5, ViDiT-Qβs W4A8 outputs show washed-out textures in elements like launch towers. In contrast, our method consistently recovers these subtle features, demonstrating robust performance. The advantage becomes more evident in more challenging W4A4 settings, where conventional methods fail completely, generating either incoherent noise patterns or severely distorted outputs. Our method maintains remarkable visual coherence with marginal quality degradation compared to the BF16 baseline.
@@ 4.3 Ablation Study
BGR and ARQ.β We validate the effectiveness of BGR and ARQ in Tab.Β 3. The full model with both BGR and ARQ achieves optimal performance across all metrics, demonstrating their synergistic effects. Removing only BGR leads to significant drops in VBench scores, indicating the critical role of progressively narrowing the weight quantization bound to preserve perceptual fidelity. Conversely, disabling ARQ components results in comparable declines, highlighting the importance of dynamically adjusting quantization scales during inference. Both BGR and ARQ maintain their effectiveness irrespective of -GBS.
| Method | -GBS | STP | ITP | ABS | SBA |
@@
| Methods | Bit Width (W/A) | Aesthetic Quality | Imaging Quality | Subject Consist. | Motion Smooth. | BG. Consist. | |
| BGR | ARQ | ||||||
| β | W4A6 | 58.15 | 58.68 | 98.04 | 98.49 | 98.21 | |
| β | W4A6 | 57.85 | 57.72 | 98.23 | 97.86 | 98.10 | |
| β | β | W4A6 | 60.46 | 61.93 | 98.91 | 98.95 | 98.40 |
| β | W4A4 | 53.95 | 52.67 | 97.92 | 98.71 | 97.89 | |
| β | W4A4 | 43.26 | 58.31 | 95.36 | 96.08 | 97.35 | |
| β | β | W4A4 | 59.57 | 58.93 | 98.67 | 99.00 | 98.47 |
Comparison with other mixed-precision strategies.β In the experiment, we set bit and bit in Eq.Β 11. Our method achieves an average bit width of 6 across 50 timesteps. We compare the proposed -Guided Bit Switching with other fixed-pattern mixed-precision strategies. Specifically, we compare with four representative approaches: (1) Static Temporal Partitioning (STP): allocate 4-bit precision for the first 25 timesteps and 8-bit precision for the subsequent 25 timesteps; (2) Inverse Temporal Partitioning (ITP): reverse the bitwidth allocation order of STP with 8-bit initially and 4-bit later; (3) Alternating Bitwidth Switching (ABS): dynamically alternate between 4-bit and 8-bit precision at each timestep; (4) Stochastic Bitwidth Allocation (SBA): randomly assign 4-bit or 8-bit precision to each timestep while maintaing the average bit number 25. As shown in Tab.Β 2, -Guided Bit Switching surpasses other static mixed-precision strategies in imaging quality. This error-driven dynamic decision mechanism effectively resolves the challenge of bit allocation in conventional mixed-precision strategies.
Threshold of -Guided Bit Switching.β In our mixed-precision strategy, the threshold determines when to switch bit-width during inference. Empirically, a lower leads to more time steps using high-bit quantization, improving visual quality at the cost of increased bit-width. This presents a trade-off between performance and resource usage. Thus, we conduct a series of experiments with different thresholds. As shown in Fig.Β 6, our experiments demonstrate that as increases, the average bit-width decreases, leading to a gradual reduction in imaging quality. Notably, a key advantage of -GBS over existing methods is its smooth and continuous bit-width adaptation: when , the model consistently uses W4A8; when , it reduces to W4A4. Critically, for intermediate values, -GBS dynamically interpolates between 4-bit and 8-bit at the activation level, ensuring a continuous precision transition based on input characteristics. This smooth adaptation is in stark contrast to conventional approaches, which typically switch abruptly between discrete bit-widths, causing instability in output quality.
| Method | Bit-width (W/A) | Aesthetic Quality | Imaging Quality | BG. Consist. | Subject. Consist. | Motion Smooth. | Speedup |
| -Quant+TeaCache | W4A8 | 61.39 | 62.96 | 98.24 | 98.73 | 98.75 | 4.01 |
@@ 4.4 Memory and Latency
-Quant achieves significant improvements in both memory efficiency and inference speed. As shown in Tab.Β LABEL:tab:storage, compared to the BF16 baseline (16/16 bit-width), our W4A8 configuration reduces memory usage by 3.68 while accelerating latency by 1.75. More aggressive quantization (W4A4) maintains the same memory savings but further boosts latency speedup to 2.12. Notably, W4A6 means roughly half the denoising steps use W4A8 and the other half use W4A4. This allows direct use of the widely adopted W4A4 and W4A8 GEMM kernels, eliminating the need to design dedicated mixed-precision kernels.
@@ 4.5 Integrate with Cache Mechanism
To demonstrate the orthogonal compatibility of -Quant with other compression paradigms, we integrate it with one of the SOTA cache methods TeaCacheΒ (Liu et al., 2024a). As illustrated in Tab.Β LABEL:tab:teacache, -Quant maintains its compression efficiency alongside TeaCacheΒ (Liu et al., 2024a), achieving additive speedup (up to 4.85) without degrading key video generation metrics. This validates the practicality of -Quant in resource constrained scenarios where multiple optimization dimensions could be jointly addressed.